OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. y – 2 = 3(x – 5) y + 8 = 3(x + 5) y + 5 = 3(x + 2) y – 5 = 3(x – 2)

OpenStudy (lgbasallote):

if i told you the point-slope frm is \[y - y_1 = m (x - x_1)\] where m = slope x1 = x coordinate y1 = y coordinate will that help you @evergirl ?

OpenStudy (anonymous):

a little not really

OpenStudy (lgbasallote):

what is your slope here?

OpenStudy (anonymous):

idk

OpenStudy (lgbasallote):

hint: look at your question

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

it is stated right there

OpenStudy (lgbasallote):

yes so m =3

OpenStudy (lgbasallote):

now what is the x coordinate?

OpenStudy (anonymous):

(2, 5)

OpenStudy (lgbasallote):

the x coordinate. not the point. you know what the x-coordinate is, right?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

 got it now?

OpenStudy (anonymous):

It is clear that y – 5 = 3(x – 2) It has slope 3 and if x=2, then y=5